# Impact of the relativistic Cowling approximation on shear and interface modes of neutron stars

### v1

@@ IV.1.2 Characteristics of $s$-modes for different values of $l$
Having established a way of determining (or at least estimating) the damping times of $s$-modes, we will now, for completeness, investigate the $s$-modes for higher polar order $l>2$. It is well-known that the frequencies of the $f$- and $p$-modes depend strongly on $l$; their frequencies increase with increasing $l$. For the $s$-modes, however, we find that the frequencies are nearly independent of $l$; this has, to the best of our knowledge, been pointed out only once by Vavoulidis et al. [74]. We show the spectra for $2\leq l\leq 7$ of the NS model with $\epsilon_{c}=0.907\times 10^{15}\textrm{g}/\textrm{cm}^{3}$ in Fig. 5. The first eight $s$-modes are clearly visible; their frequency is virtually insensitive to the value of $l$ (in contrast to the $f$-mode frequency). We show the frequencies exemplarily for three of the $s$-modes, i.e., $s_{1}$-, $s_{2}$-, and $s_{5}$-modes, in Table 5; they are not identical for different values of $l$ but vary only by tiny amounts. Except for the $s_{1}$-mode, whose frequency shows an opposite behavior up to $l=5$, the frequencies very slowly increase with increasing value of $l$. We do not show a graph with (the radial component of) the eigenfunctions but we find that those are also effectively insensitive to the value of $l$ like the frequencies.
@@
The damping times, however, change drastically with increasing value of $l$. Making use of the features of the associated Breit-Wigner resonances, we are able to estimate the damping times of the $s$-modes for $l=2$ and $l=3$; we show those values in Table 6. For higher values of $l$ (i.e., $l\geq 4$), the damping times are so large that a frequency resolution of $\Delta\omega=4\pi\times 10^{-14}$ is not sufficient to resolve the different poles and zeroes. We can, therefore, only provide the lower bound $\tau>8000\times 10^{6}\,\textrm{s}\approx 250\,\textrm{yrs}$ for the damping times of higher order $s$-modes. Looking at Table 6, we notice that the damping times of the $s$-modes with $l=3$ are very roughly three orders of magnitude larger than those for $l=2$ (our results indicate that only a slightly larger frequency resolution could reveal the features for the $l=3$ $s_{1}$-mode, suggesting that a similar factor also holds for this mode). Not being able to resolve the features for the $l=4$ $s_{5}$-mode, we expect a similar increase in damping times for higher values of $l$, too.
@@
2 661.165 1075.525 2058.492
3 661.081 1075.607 2058.747
4 661.001 1075.723 2059.092
5 660.953 1075.880 2059.525
6 660.967 1076.085 2060.047
7 661.079 1076.346 2060.658
Table 6: Damping times of the $s$-modes reported in Table 5 in multiples of $10^{6}\,$s; they increase quickly with increasing $l$. We are unable to determine numerically the damping times of $s$-modes for $l\geq 4$ due to finite resolution, but can only provide a lower bound of $\tau\gtrsim 10^{10}\,$s for these.
$l$ $s_{1}$ $s_{2}$ $s_{5}$
2 19.1 5.15 0.043
3 $>8000$ 10240 21.7
@@ IV.2 Interface Modes
In Fig. 6, we show the low-frequency domain of the spectrum for the NS model with $\epsilon_{c}=0.536\times 10^{15}\textrm{g}/\textrm{cm}^{3}$. Contrary to the shear modes (cf. Sec. IV.1), whose frequencies displayed perfect agreement between the Cowling results and those in full General Relativity, we observe differing frequencies for the $i$-modes. We explicitly show the frequencies of the two previously considered models in Tables 7 and 8; other NS models show comparable behavior.
@@
Figure 7: Comparison of the eigenfunctions of the tangential displacement $\tilde{V}$ of the two interface modes of the NS model with $\epsilon_{c}=1.626\times 10^{15}\,\textrm{g}/\textrm{cm}^{3}$ and $\alpha=1.0$; in order to distinguish the modes, we use light colors for the $i_{2}$-mode. Besides the difference in frequency (cf. Table 8) between the full GR solution and the Cowling approximation, we also observe a difference in the eigenfunctions: for the $i_{1}$-mode the discontinuity is stronger in full GR, while it is larger in the Cowling approximation for the $i_{2}$-mode.
@@
The discrepancy of the frequencies as well as eigenfunctions of the $i$-modes is somewhat unexpected. As they are modes that couple only very weakly to the spacetime and hardly radiate gravitational waves, we would expect them to be virtually unaffected by the Cowling approximation in the same way that we have observed it for the $s$-modes (cf. Sec. IV.1). In order to resolve this observation, we have first compared the background solution of our two independent TOV solvers; they yield nearly identical NS models whose radii, at which the surface and the crustal boundaries are located, differ only on the order of centimeters. While this may have an impact on the resulting frequencies, we are not convinced that this can explain the observed frequency discrepancies of several percent.
We also investigate the damping time of the $i$-modes. For both interface modes of the NS model with $\epsilon_{c}=1.626\times 10^{15}\textrm{g}/\textrm{cm}^{3}$ and $\alpha=1.0$ we are unable to resolve the individual poles and zeroes that are required for the method detailed in App. B. To be precise, at a frequency resolution of $\Delta\omega=2\pi\times 10^{-9}\,$z the necessary features cannot be resolved; at a 10-fold higher frequency resolution $\Delta\omega=2\pi\times 10^{-10}\,$z numerical inaccuracies are visible in the graph which prevent an accurate determination of the phase of the Breit-Wigner resonance. We can, therefore, only conclude that the damping time $\tau$ of both interface modes is $\tau>3.18\times 10^{11}\,\textrm{ms}\approx 10\,\textrm{yrs}$ (since $\beta_{n}<\Delta\omega/2=\pi\times 10^{-9}\,\textrm{Hz}$ according to our analysis in App. B).
@@ V Conclusions and Outlook
We construct relativistic neutron star models that possess an elastic, albeit unstrained, crust composed of spherical and cylindrical nuclei. On these background models, we calculate shear and interface modes, which are due to the presence of the elastic crust and its boundaries, respectively, and compare their characteristics in full General Relativity as well as the relativistic Cowling approximation (in which spacetime perturbations are neglected). We find that the frequency as well as the eigenfunctions of the shear modes are virtually unaffected by the Cowling approximation and the marginal differences between the obtained frequencies can be attributed to numerical inaccuracies in the respective implementations of the independently developed codes. Additionally, we observe that the frequency of the shear modes is only marginally affected by the polar quantum number $l$. In contrast, we find that the frequencies of the interface modes are altered by a few percent by the Cowling approximation; their eigenfunctions differ somewhat from each other, too. This result is somewhat unexpected and the discrepancies in the frequencies and eigenfunctions are considerably larger than the marginal differences that we detected between the solutions of our respective TOV solvers. While we do not commit ourselves to stating which of our two numerical codes produces the “correct” results, we simply advise caution in the calculation of $i$-modes as even tiny differences in the constructed background model may lead to deviations in the results that are larger than perhaps initially expected.
@@
Further, we have extended a previously published method that utilizes the features of the Breit-Wigner resonances, namely poles and zeroes in its phase, which are associated with each mode, in order to determine the damping times of slowly damped modes. We demonstrate the accuracy of this method and how to apply it on the $f$- and $p$-modes, whose damping times can be accurately determined also by minimizing the error function $|A_{\textrm{in}}|$ in the complex plane. Sometimes, not all poles or zeroes can be determined or it could be the case that their location is beyond the validity of the assumed linearity approximation. For these cases, where only a subset of the features is accurately known, we show how the damping time of a slowly damped mode can still be accurately calculated or at least be robustly constrained. Employing these extensions, we accurately determine the damping times of shear modes; we find that the first few overtones have considerably shorter damping times than the fundamental shear mode. When increasing the polar order $l$, the damping time of the shear modes quickly increases, too. We also attempt to estimate the damping time of the interface modes; however, these are so long that we can only provide a lower bound of $\approx 10\,$yrs for them.
@@ Acknowledgements.
C.K. warmly thanks the Interdisciplinary Theoretical and Mathematical Science Program (iTHEMS) of RIKEN for their kind hospitality and support by the iTHEMS working group “GW-EOS”, where this work has been completed. This work is supported in part by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Numbers JP23K20848 and JP24KF0090, and by FY2023 RIKEN Incentive Research Project.
@@ Appendix B Constraining damping times of long-lived QNMs
For simplicity, we keep the notation used in Ref. [72] and repeat the most important details: The method is based on the ratio $\kappa(\omega):=A_{\textrm{out}}(\omega)/A_{\textrm{in}}(\omega)$ of the amplitudes of the outgoing wave $A_{\textrm{out}}$ and the ingoing wave $A_{\textrm{in}}$ at infinity. The relativistic system has a QNM with frequency $\omega_{n}=\alpha_{n}+i\beta_{n}$ if $A_{\textrm{in}}(\omega_{n})=0$; at the same time, we have $A_{\textrm{out}}(\overline{\omega}_{n})=0$.
@@ B.1 Testing the method
[kHz] [ms]
complex iteration 5.837954473 5086.39 -5.620
estimate 1 5.837954474 5086.43 -5.619
estimate 2 5086.44 -5.618
@@
[kHz] [ms]
complex iteration 7.808690352 438437 -1.437
estimate 1 7.808690360 445753 -1.368
estimate 2 424728 -1.370
@@
We now turn to the $f$-mode which can be easily iterated for in the complex plane. On the real axis, we find two poles and one zero in the vicinity of the $f$-mode frequency (which is 1.8296183131.8296183131.829618313\,1.829618313kHz, i.e., marginally smaller than $Z_{2}$), but somewhat surprisingly, we are unable to locate the second zero. Scanning the complex plane, we find that $\mathcal{G}\approx 545>0$, which implies that $Z_{1}<P_{1}$ (cf. Eq. (46)). In fact, we should find $Z_{1}$ roughly at 1.31.31.3\,1.3kHz; however, the linearity approximation no longer holds at this distance which explains why we are not able to locate a (reliable) corresponding $Z_{1}$.
@@
$P_{1}/2\pi$ 1.829135836
$Z_{2}/2\pi$ 1.829618756
$P_{2}/2\pi$ 1.830102563
Using the correct value of $\alpha_{n}$ (e.g., by finding the minimum of the Breit-Wigner resonance for real-values $\omega$), we can employ $P_{1}$ and $P_{2}$ in Eq. (40) and find $\tau=329.26603\,$ms which is only marginally different from the correct value $\tau=329.26612\,$ms. Alternatively, without employing the correct value of $\alpha_{n}$, we can use Eq. (53) and find $329.265\,\textrm{ms}<\tau\leq 465.65\,\textrm{ms}$; secondly, employing Eq. (51) we find $273.023\,\textrm{ms}<\tau<329.568\,\textrm{ms}$. These two intervals only have a tiny overlap region which contains the true value of $\tau$.
Analyzing the $f$- and $p$-modes of different NS models, we make very similar findings in the accuracy of the estimates of frequency and damping time. Overall, we can confirm that this method works very well and even if the features of the Breit-Wigner resonance are incomplete, a quite reliable value for the damping time can be extracted based on the presented inequalities.

### v2

@@ V.1.2 Characteristics of $s$-modes for different values of $l$
Having established a way of determining (or at least estimating) the damping times of $s$-modes, we will now, for completeness, investigate the $s$-modes for higher polar order $l>2$. It is well-known that the frequencies of the $f$- and $p$-modes depend strongly on $l$; their frequencies increase with increasing $l$. For the $s$-modes, however, we find that the frequencies are nearly independent of $l$; this has, to the best of our knowledge, been pointed out only once by Vavoulidis et al. [78]. We show the spectra for $2\leq l\leq 7$ of the NS model with $\epsilon_{c}=0.907\times 10^{15}\textrm{g}/\textrm{cm}^{3}$ in Fig. 5. The first eight $s$-modes are clearly visible; their frequency is virtually insensitive to the value of $l$ (in contrast to the $f$-mode frequency). We show the frequencies exemplarily for three of the $s$-modes, i.e., $s_{1}$-, $s_{2}$-, and $s_{5}$-modes, in Table 5; they are not identical for different values of $l$ but vary only by tiny amounts. Except for the $s_{1}$-mode, whose frequency shows an opposite behavior up to $l=5$, the frequencies very slowly increase with increasing value of $l$.
@@
The damping times, however, change drastically with increasing value of $l$. Making use of the features of the associated amplitude ratios, we are able to estimate the damping times of the $s$-modes for $l=2$ and $l=3$; we show those values along with estimates from the quadrupole formulas in Table 6. For higher values of $l$ (i.e., $l\geq 4$), the damping times are so large that a frequency resolution of $\Delta\omega=4\pi\times 10^{-14}$ is not sufficient to resolve the different poles and zeroes. We can, therefore, only provide the lower bound $\tau>8000\times 10^{6}\,\textrm{s}\approx 250\,\textrm{yrs}$ for the damping times of higher order $s$-modes. The damping times obtained by means of the quadrupole formulas, again, systematically underestimate the values obtained by analysing the phase of the amplitude ratio; however, in contrast to the results shown in Tab. 4, the values from the relativistic modification R/RQF deviate less from the expected value than those from the N/SQF expression.
Looking at Table 6, we notice that the damping times of the $s$-modes with $l=3$ are very roughly three orders of magnitude larger than those for $l=2$ (our results indicate that only a slightly larger frequency resolution could reveal the features for the $l=3$ $s_{1}$-mode, suggesting that a similar factor also holds for this mode). Not being able to resolve the features for the $l=4$ $s_{5}$-mode, we expect a similar increase in damping times for higher values of $l$, too. We note that the eigenfunctions of the $s$-modes of different $l$ are qualitatively quite similar; this is particularly true within the solid part of the crust where those functions are significantly different from zero and, hence, provide their main contribution to the integrals in the expressions (30) and (31). However, the amplification factors vary slightly between different modes and also the eigenfunctions of the perturbed densities (both rest-mass and energy) display some larger mismatches. If those eigenfunctions were “identical” (up to certain amplification factors), the damping time of an $l>2$ $s$-mode could be calculated from the damping time of the $l=2$ $s$-mode. Our current data set is not large enough to provide a reliable estimate for these amplification factors at this time.
@@
2 654.887 1068.20 2056.48
3 654.811 1068.29 2056.72
4 654.745 1068.41 2057.06
5 654.716 1068.58 2057.49
6 654.757 1068.80 2058.01
7 654.904 1069.08 2058.61
Table 6: Damping times of the $s$-modes reported in Table 5 in multiples of $10^{6}\,$s; they increase quickly with increasing $l$. We are unable to determine numerically the damping times of $s$-modes for $l\geq 4$ due to finite resolution, but can only provide a lower bound of $\tau\gtrsim 10^{10}\,$s for these. PAR denotes the damping times estimated via the phase of the amplitude ratio, while N/SQF and R/RQF are estimates employing variants of the quadrupole formula.
Damping time $[10^{6}\,\textrm{s}]$
$l$ Method $s_{1}$ $s_{2}$ $s_{5}$
2 PAR 15.2 2.29 0.0099
N/SQF 11.0 3.20 0.0290
R/RQF 14.8 4.22 0.0370
3 PAR $>8000$ $>8000$ 6.46
N/SQF 30000 4836 10.7
R/RQF 40400 6399 14.1
@@ V.2 Interface Modes
In Fig. 6, we show the low-frequency domain of the spectrum for the NS model with $\epsilon_{c}=0.536\times 10^{15}\textrm{g}/\textrm{cm}^{3}$. Contrary to the shear modes (cf. Sec. V.1), whose frequencies displayed perfect agreement between the Cowling results and those in full General Relativity, we observe differing frequencies for the $i$-modes. We explicitly show the frequencies of the two previously considered models in Tables 7 and 8; other NS models show comparable behavior.
@@
Figure 7: Comparison of the eigenfunctions of the tangential displacement $\tilde{V}$ of the two interface modes of the NS model with $\epsilon_{c}=1.626\times 10^{15}\,\textrm{g}/\textrm{cm}^{3}$ and $\alpha=1.0$; We show the $i_{1}$-mode in red and the $i_{2}$-mode in blue; full GR and Cowling results are distinguished by solid and dashed lines, respectively. Besides the difference in frequency (cf. Table 8) between the full GR solution and the Cowling approximation, we also observe a difference in the eigenfunctions: for the $i_{1}$-mode the discontinuity is stronger in full GR, while it is larger in the Cowling approximation for the $i_{2}$-mode.
@@
The discrepancy of the frequencies as well as eigenfunctions of the $i$-modes is somewhat unexpected. As they are modes that couple only very weakly to the spacetime and hardly radiate gravitational waves, we would expect them to be virtually unaffected by the Cowling approximation in the same way that we have observed it for the $s$-modes (cf. Sec. V.1). In order to resolve this observation, we have first compared the background solution of our two independent TOV solvers; they yield nearly identical NS models whose radii, at which the surface and the crustal boundaries are located, differ only on the order of centimeters. While this may have an impact on the resulting frequencies, we are not convinced that this can explain the observed frequency discrepancies of several percent.
We also investigate the damping time of the $i$-modes. For both interface modes of the NS model with $\epsilon_{c}=1.626\times 10^{15}\textrm{g}/\textrm{cm}^{3}$ and $\alpha=1.0$ we are unable to resolve the individual poles and zeroes that are required for the method detailed in App. B. To be precise, at a frequency resolution of $\Delta\omega=2\pi\times 10^{-9}\,$z the necessary features cannot be resolved; at a 10-fold higher frequency resolution $\Delta\omega=2\pi\times 10^{-10}\,$z numerical inaccuracies are visible in the graph which prevent an accurate determination of the phase of the amplitude ratio. We can, therefore, only conclude that the damping time $\tau$ of both interface modes is $\tau>3.18\times 10^{11}\,\textrm{ms}\approx 10\,\textrm{yrs}$ (since $\beta_{n}<\Delta\omega/2=\pi\times 10^{-9}\,\textrm{Hz}$ according to our analysis in App. B).
@@ VI Conclusions and Outlook
We construct relativistic neutron star models that possess an elastic, albeit unstrained, crust composed of spherical and cylindrical nuclei. On these background models, we calculate shear and interface modes, which are due to the presence of the elastic crust and its boundaries, respectively, and compare their characteristics in full General Relativity as well as the relativistic Cowling approximation (in which spacetime perturbations are neglected). We find that the frequency as well as the eigenfunctions of the shear modes are virtually unaffected by the Cowling approximation and the marginal differences between the obtained frequencies can be attributed to numerical inaccuracies in the respective implementations of the independently developed codes. Additionally, we observe that the frequency of the shear modes is only marginally affected by the polar quantum number $l$. In contrast, we find that the frequencies of the interface modes are considerably by the Cowling approximation; their eigenfunctions differ from each other, too. This result is somewhat unexpected and the discrepancies in the frequencies and eigenfunctions are considerably larger than the marginal differences that we detected between the solutions of our respective TOV solvers. While we do not commit ourselves to stating which of our two numerical codes produces the “correct” results, we simply advise caution in the calculation of $i$-modes as even tiny differences in the constructed background model may lead to deviations in the results that are larger than perhaps initially expected.
@@
Further, we have extended a previously published method that utilizes the features of the phase of the amplitude ratio, namely poles and zeroes in its phase, which are associated with each mode, in order to determine the damping times of slowly damped modes. We demonstrate the accuracy of this method and how to apply it on the $f$- and $p$-modes, whose damping times can be accurately determined also by minimizing the error function $|A_{\textrm{in}}|$ in the complex plane. Sometimes, not all poles or zeroes can be determined or it could be the case that their location is beyond the validity of the assumed linearity approximation. For these cases, where only a subset of the features is accurately known, we show how the damping time of a slowly damped mode can still be accurately calculated or at least be robustly constrained. Employing these extensions, we accurately determine the damping times of shear modes; we find that the first few overtones have considerably shorter damping times than the fundamental shear mode. When increasing the polar order $l$, the damping time of the shear modes quickly increases, too. We also employ the quadrupole formula for the damping times and find that it yields good order-of-magnitude estimates. We also attempt to estimate the damping time of the interface modes; however, these are so long that we can only provide a lower bound of $\approx 10\,$yrs for them.
@@ Acknowledgements.
We thank the anonymous referee for very constructive comments that helped improve the content of this manuscript. C.K. warmly thanks the Interdisciplinary Theoretical and Mathematical Science Program (iTHEMS) of RIKEN for their kind hospitality and support by the iTHEMS working group “GW-EOS”, where this work has been completed. This work is supported in part by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Numbers JP23K20848 and JP24KF0090, and by FY2023 RIKEN Incentive Research Project.
@@ Appendix B Constraining damping times of long-lived QNMs
For simplicity, we keep the notation used in Ref. [71] (however, we capitalize the variables for the poles and zeroes to avoid confusion with $p_{1}$- and $p_{2}$-modes) and repeat the most important details: The method is based on the ratio $\kappa(\omega):=A_{\textrm{out}}(\omega)/A_{\textrm{in}}(\omega)$ of the amplitudes of the outgoing wave $A_{\textrm{out}}$ and the ingoing wave $A_{\textrm{in}}$ at infinity. The relativistic system has a QNM with frequency $\omega_{n}=\alpha_{n}+i\beta_{n}$ if $A_{\textrm{in}}(\omega_{n})=0$; at the same time, we have $A_{\textrm{out}}(\overline{\omega}_{n})=0$.
@@ B.1 Testing the method
[kHz] [s]
complex iteration 5.837974945 5.07694 -5.620
estimate 1 5.837974945 5.07693 -5.621
estimate 2 5.07870 -5.620
@@
[kHz] [s]
complex iteration 7.808828031 438.268 -1.436
estimate 1 7.808828023 431.626 -1.508
estimate 2 453.390 -1.510
@@
We now turn to the $f$-mode which can be easily iterated for in the complex plane. On the real axis, we find two poles and one zero in the vicinity of the $f$-mode frequency (which is 1.8296263271.8296263271.829626327\,1.829626327kHz, i.e., marginally smaller than $Z_{2}$), but somewhat surprisingly, we are unable to locate the second zero. Scanning the complex plane, we find that $\mathcal{G}\approx 545>0$, which implies that $Z_{1}<P_{1}$ (cf. Eq. (55)). In fact, we should find $Z_{1}$ roughly at 1.31.31.3\,1.3kHz; however, the linearity approximation no longer holds at this distance which explains why we are not able to locate a (reliable) corresponding $Z_{1}$.
@@
$P_{1}/2\pi$ 1.829143927
$Z_{2}/2\pi$ 1.829626828
$P_{2}/2\pi$ 1.830110615
Using the correct value of $\alpha_{n}$ (e.g., by finding the minimum of the Breit-Wigner resonance for real-values $\omega$), we can employ $P_{1}$ and $P_{2}$ in Eq. (49) and find $\tau=329.27942\,$ms which is only marginally different from the correct value $\tau=329.27944\,$ms. Alternatively, without employing the correct value of $\alpha_{n}$, we can use Eq. (62) and find $329.279\,\textrm{ms}<\tau\leq 465.67\,\textrm{ms}$; secondly, employing Eq. (60) we find $273.033\,\textrm{ms}<\tau<329.581\,\textrm{ms}$. These two intervals only have a tiny overlap region which contains the true value of $\tau$.
Analyzing the $f$- and $p$-modes of different NS models, we make very similar findings in the accuracy of the estimates of frequency and damping time. Overall, we can confirm that this method works very well, and even if the features of the phase of the amplitude ratio are incomplete, a quite reliable value for the damping time can be extracted based on the presented inequalities.